OpenStudy (anonymous):

what is the inverse of the matrix??? [-8 5; 8 -3]

OpenStudy (anonymous):

1/-16(8 8/5 3)

OpenStudy (anonymous):

i think it might be negate a11b22 and switch b11a22

OpenStudy (anonymous):

NOPE this one above is not correct you need 1/determinant

OpenStudy (anonymous):

bnut is correct, dont forget to divide by the det of the original matrix.

OpenStudy (anonymous):

Yeah...So the determinant is -8(-3)-40=-16 So you have: \[A^{-1}=\left[\begin{matrix}\frac{3}{16} & \frac{5}{16} \\ \frac{1}{2} & \frac{1}{2}\end{matrix}\right]\]

OpenStudy (anonymous):

I thought it in my head but just forgot to type it lol

OpenStudy (anonymous):

i always forget which one to negate and which two to flip. you could alway to the adjoint